OpenStudy (anonymous):

(Newton's Law of Motion) A billy goat running at 5 m/s butts his head against a tree. If his head has a mass of 4 kg and it comes to a stop in 5 millimeters, find the average force of the impact. What do you think will happen to the goat?

OpenStudy (anonymous):

5*5=2*a*.005 a=2500 m/s^2 f=4*2500 f=10000 N I am sorry for billy!! but where is he now ?